OpenStudy (anonymous):

Simplify. 2 + 4 • (6 – 3)

OpenStudy (yanasidlinskiy):

the answer is 9. use pemdas

OpenStudy (anonymous):

i have another question

OpenStudy (yanasidlinskiy):

what is it...

OpenStudy (anonymous):

wait about the 1st question thats not true

OpenStudy (anonymous):

these are my options A. 18 B. 14 C. 23 D. 33

OpenStudy (anonymous):

and its ok we all make mistakes

OpenStudy (yanasidlinskiy):

wait...my bad it is actually 18.. Sorry for the mistake i read it wrong:) but its 18

OpenStudy (anonymous):

I got 14

OpenStudy (anonymous):

oooooooo its ok i thought u were mad at me

OpenStudy (anonymous):

Yes it's 14.

OpenStudy (anonymous):

2+4*(6-3) 2+4*(3) 2+(12) =14

OpenStudy (anonymous):

oooooooooooooooooooooooo ok i have another question

OpenStudy (anonymous):

@selenatorssmiler4life is right. The answer is 14. You have to multiply 4*3 BEFORE you add the 2 :-D

OpenStudy (anonymous):

Where's the next problem ?

OpenStudy (anonymous):

A. 4 B. 3 C. 6 D. 16

OpenStudy (anonymous):

ooo sorry

OpenStudy (anonymous):

2+4 X(6-3) 2+4=6 6-3=3 6x3=18

OpenStudy (anonymous):

simplify the problem 18-2*3 __________ 12/6+1

OpenStudy (anonymous):

these r my options A. 4 B. 3 C. 6 D. 16

OpenStudy (anonymous):

4

OpenStudy (anonymous):

6a-2b+5 Evaluate the expression when a = 2 and b = 4.

OpenStudy (anonymous):

my options A. 43 B. 45 C. 1 D. 9

OpenStudy (anonymous):

http://sketchtoy.com/59463839

OpenStudy (anonymous):

The answer is A 18-(2*3) = 18-6= 12 ------------------- (12/6)+1 = 2+3= 3 12 --- =4 3

OpenStudy (anonymous):

Answer is D=9

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

ooo ok

OpenStudy (yanasidlinskiy):

i got d. 9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

next question

OpenStudy (yanasidlinskiy):

6(2)-2(4)+5 12-8 +5 17-8 9.....np:)

OpenStudy (anonymous):

(1 pt) Which is a variable expression for the word phrase? the product of 3 and a number decreased by 7 A. 3n – 7 B. 3n + 7 C. 7n – 3 D. 7 – 3n

OpenStudy (yanasidlinskiy):

my opinion would be d. you might want t oget another opinion...

OpenStudy (anonymous):

ok then d it is

OpenStudy (anonymous):

The answer is actually A= 3n-7

OpenStudy (anonymous):

it would be D if it said 7 decreased by the product of 3and a number.

OpenStudy (yanasidlinskiy):

i was actually thinking that it might be 3n-7 but i don't know.

OpenStudy (anonymous):

ok then a it is

OpenStudy (anonymous):

Select the mathematical sentence that is true. A. 2 • (8 – 5) + 3 > 7 + 3 • 2 B. 2 • (8 – 5) + 3 < 7 + 3 • 2 C. 2 • (8 – 5) + 3 = 7 + 3 • 2

OpenStudy (anonymous):

anymore questions?

OpenStudy (anonymous):

Select the mathematical sentence that is true. A. 2 • (8 – 5) + 3 > 7 + 3 • 2 B. 2 • (8 – 5) + 3 < 7 + 3 • 2 C. 2 • (8 – 5) + 3 = 7 + 3 • 2

OpenStudy (anonymous):

can u see it

OpenStudy (anonymous):

The answer is B.

OpenStudy (anonymous):

ooo ok

OpenStudy (anonymous):

3 more

OpenStudy (anonymous):

Bring it!

OpenStudy (anonymous):

Find the solution set for the given inequality if the replacement set is {1, 2, 3, 4, 5}. 5x – 2 ≥ 8 A. {3, 4, 5} B. {2, 3, 4, 5} C. {1, 2} D. {1}

OpenStudy (anonymous):

The answer is B :D

OpenStudy (anonymous):

Determine which value is a solution to the open sentence. 6 + 2x = 24 A. 9 B. 16 C. 15 D. 6

OpenStudy (anonymous):

The answer is A.

OpenStudy (anonymous):

Which is a word phrase for the variable expression? 4(n – 2) A. four times the difference of a number n and 2 B. four times the product of a number n and 2 C. four more than the difference of a number n and 2 D. two less than four times a number n

OpenStudy (anonymous):

The answer is A.

OpenStudy (anonymous):

ooo ok thanks for evrything and for thinking somthing i cant do hahahahah

OpenStudy (anonymous):

No prob! Have a good one :-D